OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. 4x – 2y – z = –5 x – 3y + 2z = 3 3x + y – 2z = –5

OpenStudy (anonymous):

damn, you look hot. add 2nd and 3rd equation: x - 3y + 2z + 3x + y - 2z = -2 4x - 2y = -2 substitute this into the first equation: -2 - z = -5 therefore z = 3

OpenStudy (anonymous):

therefore: 3x + y = 1 4x - 2y = -2 multiply first equation by 2: 6x + 2y = 2 add this to second equation: 10x = 0 therefore x = 0, therefore y = 1

OpenStudy (anonymous):

ur missing one

OpenStudy (anonymous):

(0, –1, –3) (0, –1, 3) (0, 1, 3) (0, 1, –3)

OpenStudy (anonymous):

no, i'm not. i said z = 3 it's (0, 1, 3)

OpenStudy (anonymous):

sorry hun i did not see it

OpenStudy (anonymous):

im falling asleep

OpenStudy (anonymous):

that's ok :)